I am confused on the idea os "N-th" root & "cubic" root. I barely understand square root. Is there an easy way to explain this?
Answer by stanbon(75887)

I am confused on the idea of "N-th" root & "cubic" root. I barely understand square root. Is there an easy way to explain this?
-----
square root of 25 = 5 because 25 = 5*5
3rd root of 125 = 5 because 125 = 5*5*5
4th root of 625 = 5 because 625 = 5*5*5*5
-----
nth root of "x" is the number which when used
as a factor n time gives a product of "x".
=================================
cubic root means 3rd root
=================================
Cheers,
Stan H.